# Are EU low-carbon structural funds efficient in reducing emissions?

### v1

@@ 4.1 Effect on emissions trend
| Notes: NUTS 2 cluster robust standard errors are in parentheses. Significance level: ∗p0.1; ∗∗p0.05; ∗∗∗p0.01. | ||||||
Therefore, for developed regions, the reference group, the effect of expenditures in both periods indicates a negative and significant impact in all specifications, except in model 3, where the effect is not significant for both periods. Similarly, the effect of the interaction term for transition regions is negative in all specifications, except in model 2, being non-significant for expenditures associated with the period 2007-2013 and significant for expenditures in the period 2014-2020. This implies that the net impact for transition regions is negative, meaning that, for both transition and developed regions, the long-term effect of expenditures is expected to result in a decrease in emissions, both per capita and in relation to GDP.
In contrast, a significant positive effect is observed for less developed regions, reversing the trend seen in the other development groups. This is because the estimates of the expenditures interaction terms in developing regions are not only positive and significant in all specifications but also cancel any negative effect estimated for the reference group. While explaining the mechanism behind this perverse outcome is far from the scope of this paper, this could indicate that, in less developed regions, expenditures have effectively driven economic development, albeit possibly at the cost of an increase in emissions in the long term. It is important to note that as these regions are developing, the increase in emissions resulting from factors like industrialisation, which is often carbon-intensive, cannot be ignored.
Regarding the control variables, GDP is positively related to GHGEpc and GHGEi, implying that economic growth can lead to an increase in per capita emissions in the long term and to a more efficient economy in terms of emissions relative to GDP.
The EPS dummy has a small but positive impact on per capita emissions and negative emission intensity. Although this positive impact is unexpected, it may also indicate that the effect of more stringent policies may be observed thereafter. Indeed, Naqvi (2021) analysed the cumulative effects of EPS dummy on different emission components, showing that while CH4 and N2O emissions responded strongly to policies and declined continuously across regions, CO2 emissions declined two years after implementing the policies. In the Appendix, we show that the EPS dummy has a negative impact on CH4 and N2O while positive for CO2, the latter being the main component of the total GHG emissions used in this paper (Tables A4 to A6).
The effect of the average political ideology is not robust; while it is found positive for GHGEpc, except in model 3, for GHGEi, it is positive and not significant in model 5 and negative and significant in model 6. These mixed results connect to the complex relationship between the effectiveness of sustainable policies and political ideologies. In this respect, for instance, Garmann (2014) finds that in 19 OECD countries from 1992 to 2008, right-wing governments were associated with fewer emission reductions than the centre and left-wing governments. However, he also observes that emissions were higher in governments with more participating parties, although the distinction between majority and minority governments did not significantly influence emissions.
The effect related to the gross value added by sectors reveals a differentiated impact on GHGEpc and GHGEi. For instance, the GVA of the agriculture sector, GVAA, has a negative impact on both GHGEpc and GHGEi, which suggests that the trend of the agricultural sector production contributes to a reduction of the emissions trends. While the GVA of the manufacturing industry, GVAB-E has a significant positive effect on GHGEpc and a significant negative effect on GHGEi, indicating that greater production in the manufacturing industry implies higher emissions per capita but lower emission intensity.
4.2 Effect in the emissions cycle
Table 3 presents the results of the regressions for the cyclical part of per capita emissions and emissions intensity. The results show mixed effects associated with the expenditures of both periods and the development level interacted variables. These effects are presented in reference to the most developed regions. Therefore, for developed regions, we find that the effect of expenditure in 2007-2013 is negative, although only significant in model 1. The effect of expenditure in 2014-2020 is not significant, but in specifications (2-3 and 5-6), the estimated effect is positive. These results hint that expenditure in the low-carbon economy has no robust effect on the emission cycle for the most developed regions.
| Dependent variable: | ||||||
| GHGEpc | GHGEi | |||||
| (1) | (2) | (3) | (4) | (5) | (6) | |
| Low Carbon Exp. pc | 0.081∗ | 0.027 | 0.034 | 0.068 | 0.025 | 0.022 |

### v2

@@ 4.2 Effect in the emissions cycle
| Notes: Robust standard errors clustered at the NUTS 2 level are reported in parentheses. Significance level: ∗p0.1; ∗∗p0.05; ∗∗∗p0.01 | ||||||
The estimates for per capita emissions cycles (columns 1c–3c) show that in developed regions, the direct effect of low-carbon expenditures is negative and statistically significant for the 2007–2013 programme in models 2c and 3c, suggesting that these investments helped smooth short-term fluctuations. Notably, the interaction terms for less-developed regions are also negative and statistically significant across all three models for the same programming period. This indicates that, despite the positive long-term trend effect observed in less-developed regions (see Table 2), low-carbon investments during the 2007–2013 programme contributed to dampening short-term emissions volatility. The coefficients for transition regions are generally negative but not statistically significant.
As regards emissions intensity cycles (columns 4c–6c), there is no significant direct effect of low-carbon spending for developed or transition regions. However, for less-developed regions, interaction terms for the 2007–2013 programme are again negative and statistically significant (in models 5c and 6c). As in the per capita models, the effects of the 2014–2020 programme are not statistically significant for any regional group.
Lagged GDP is negatively associated with per capita emissions cycles and positively with emissions intensity cycles, consistent with short-run adjustments in the emissions-output relationship. Sectoral GVA variables display heterogeneous effects: GVA in construction (sector F) is associated with a reduction in emissions per capita, while services (G–J) appear positively associated with emissions intensity in model 6c. Other sectors are not significantly correlated with the cyclical components of emissions.
Political ideology remains statistically not significant across specifications. However, environmental policy stringency shows a small but statistically significant positive effect in all models where it is included. This suggests that the cyclical component of emissions may co-move with the timing or enforcement of regulatory frameworks. These results are in line with recent findings for European regions, which report that emissions growth rates are highly volatile and sensitive to policy phases, particularly within the EU ETS framework (Dueñas and Mandel, 2025).
4.3 Instrumental variables estimations
To address potential endogeneity concerns related to low-carbon investments, we complement our baseline fixed-effects specification with an instrumental variables (2SLS) approach. While the use of lagged explanatory variables in the baseline models helps mitigate simultaneity bias, endogeneity may persist due to omitted variables or reverse causality. For instance, structurally more proactive regions might both attract higher low-carbon funding and implement complementary environmental measures not fully captured by our controls, potentially biasing the estimated coefficients.
We instrument per capita low-carbon expenditures from both programming periods (2007–2013 and 2014–2020) using regional structural indicators from the ARDECO database: (i) labour productivity per employee (log), (ii) consumption of fixed capital (log), (iii) hours worked per capita, and (iv) compensation of employees (log). These instruments are lagged by three years, not only to ensure temporal separation from emissions outcomes, but also based on the assumption that the allocation of funds is determined by pre-existing structural conditions. Additionally, we include regional development level dummies (TER and LDR, with DER as the reference category), which reflect the eligibility and allocation logic of the ESI funds. These variables are plausibly exogenous to short- and medium-term emissions dynamics.
First-stage results, presented in Table 4, confirm the strength of our instruments, with F-statistics exceeding conventional thresholds and statistically significant coefficients in the expected direction.
| Low Carbon Exp. pc07-13 | Low Carbon Exp. pc14-20 | |
| Labour productivity ppet-3 | -1.033*** | -0.683*** |
| (0.114) | (0.071) | |
| Hours worked per capitat-3 | 0.003*** | 0.001*** |
| (0.000) | (0.000) | |
| Compensation of Employeest-3 | -1.836*** | 0.422*** |
| (0.237) | (0.109) | |
| Consumption of fixed capitalt-3 | 1.990*** | -0.585*** |
| (0.237) | (0.110) | |
| TER | 0.970*** | 0.559*** |
| (0.117) | (0.066) | |
| LDR | 1.527*** | 1.329*** |
| (0.145) | (0.080) | |
| Time Dummies | Yes | Yes |
| F-statistic | 114.670 | 265.060 |
| Prob F | 0.000 | 0.000 |
| Observations | 1,694 | 1,687 |
| Notes: Standard errors in parentheses. Significance levels: ∗p0.1; ∗∗p0.05; ∗∗∗p0.01. TER and LDR refer to transition and less developed regions, respectively; DER is the reference group. | ||
For the second stage, we focus on the specifications that include the most comprehensive set of controls, those considering GVA disaggregated by sector. The column structure is intentionally aligned with the baseline results to facilitate direct comparison. Table 5 presents the second-stage results for both emissions performance indicators. Overall, the IV estimates reveal patterns that are qualitatively consistent with those obtained from the fixed-effects specification. To further support the validity of our instruments, we regress the second-stage residuals on the full set of instruments; the absence of significant coefficients suggests that the instruments are uncorrelated with the structural error term. Results of this test are reported in Appendix Table A4.
| Dependent variable: | ||||
| Trend | Cycle | |||
| (3t) | (6t) | (3c) | (6c) | |
| GHGEpc | GHGEi | GHGEpc | GHGEi | |
| Low Carbon Exp. pc | -0.587 | -0.224 | -0.601* | -0.144 |
| (0.393) | (0.389) | (0.345) | (0.201) | |
| Low Carbon Exp. pc | 5.692 | 4.207 | 0.460 | 0.340 |
| (3.468) | (3.847) | (0.815) | (0.846) | |
| Low Carbon Exp. pcTER | 1.399* | 1.949*** | -0.300 | 0.092 |
| (0.732) | (0.695) | (0.484) | (0.399) | |
| Low Carbon Exp. pcTER | -0.779 | 0.392 | -0.551 | -0.639 |
| (3.041) | (3.005) | (0.665) | (0.702) | |
| Low Carbon Exp. pcLDR | 1.126*** | 0.539 | 0.357 | 0.116 |
| (0.408) | (0.387) | (0.339) | (0.204) | |
| Low Carbon Exp. pcLDR | -2.350 | -2.487 | -0.556 | -0.360 |
| (3.189) | (3.462) | (0.740) | (0.772) | |
| EPSt-5 | 0.001 | -0.001 | 0.009*** | 0.011*** |
| (0.007) | (0.007) | (0.003) | (0.003) | |
| Political Ideologyt-5 | 0.006 | -0.010 | 0.003 | 0.002 |
| (0.010) | (0.013) | (0.003) | (0.003) | |
| GVAA,t-5 | 0.076** | 0.044 | -0.012 | -0.011 |
| (0.037) | (0.038) | (0.009) | (0.007) | |
| GVAB-E,t-5 | 0.213*** | -0.064 | -0.008 | -0.003 |
| (0.057) | (0.078) | (0.018) | (0.018) | |
| GVAF,t-5 | -0.110** | -0.134** | -0.046*** | 0.015 |
| (0.048) | (0.058) | (0.015) | (0.017) | |
| GVAG-J,t-5 | 0.205* | -0.129 | 0.018 | 0.078** |
| (0.113) | (0.129) | (0.035) | (0.032) | |
| GVAK-N,t-5 | -0.389*** | -0.436*** | -0.027 | 0.030 |
| (0.079) | (0.094) | (0.039) | (0.038) | |
| GVAO-U,t-5 | 0.147 | 0.303** | -0.180*** | -0.045 |
| (0.102) | (0.125) | (0.042) | (0.041) | |
| Observations | 3,408 | 3,408 | 3,408 | 3,408 |
| NUTS 2 fixed effects | Yes | Yes | Yes | Yes |
| Time fixed effects | Yes | Yes | Yes | Yes |
| Notes: Robust standard errors clustered at the NUTS 2 level are reported in parentheses. Significance level: ∗p0.1; ∗∗p0.05; ∗∗∗p0.01. Low Carbon Exp. pc07-13 and Low Carbon Exp. pc14-20 are instrumented. | ||||
For emissions per capita trends (Column 3t), low-carbon expenditures from the 2007–2013 programme are not statistically significant in developed regions, but the interaction term for less-developed regions remains positive and significant. This confirms that these regions experienced increases in long-term emissions per capita following higher investments. A key difference from the baseline is that the TER interaction becomes positive and significant, whereas it was previously negative and not significant. This suggests potential divergence in the response of transition regions under the IV approach. The 2014–2020 effects remain statistically not significant across development levels.
Regarding emissions intensity trends (Column 6t), we find a significant positive effect for the 2007-2013 expenditures in transition regions. This result departs from the baseline findings, where all interaction effects were not statistically significant.
Turning to the cyclical components, for emissions per capita (Column 3c), the 2007–2013 programme has a negative and statistically significant effect in developed regions, consistent with a smoothing effect on short-term volatility. However, the previously significant negative interaction effects for LDR are no longer significant under the IV specification. Similarly, for emissions intensity cycles (Column 6c), none of the coefficients related to expenditures are statistically significant, contrasting with the baseline results where some mitigating effects were observed.
In summary, the IV results reinforce the main conclusion of our baseline specification: low-carbon investments have different effects on emissions across different development levels. While some coefficients change in significance or sign, the overarching pattern, particularly the long-term positive association in less-developed regions, remains robust. These findings enhance confidence in the causal interpretation of the estimated effects and underscore the importance of considering regional structural differences when assessing climate policy outcomes.
4.4 Effect by type of gases
We break down emissions by specific gases, considering CO2, CH4, and N2O, to assess the differentiated impact of low-carbon investments. We adopt the same methodological approach used for total GHG emissions: decomposing the per capita and intensity time series for each gas into trend and cycle components, and estimating the effects of expenditures on these indicators following the specification in Eq.(3). Results are reported in Tables 6 and 7, corresponding to the trend and cycle components, respectively. For readability, the estimated coefficients of control variables are omitted from the tables.
Regarding the trend of CO2 emissions (upper panel of Table 6), we find no statistically significant effects of the 2007–2013 programme for the reference group (developed regions) in terms of per capita emissions. However, the interaction terms with less developed regions are consistently positive and significant across specifications. For CO2 intensity, a significant negative effect is observed in column 4 for the reference group, indicating possible efficiency gains, yet these are counteracted by a positive and significant interaction with LDR in the same column. For the 2014–2020 programme, effects on CO2 per capita and intensity are generally not significant. Nonetheless, the interaction with LDR remains positive and significant in columns 1 and 2, reinforcing the earlier programming effect pattern.
| Dependent variables by gas type | ||||||
| CO2 per capita | CO2 intensity | |||||
| Model | (1) | (2) | (3) | (4) | (5) | (6) |
| Low Carbon Exp. pc | 0.066 | -0.061 | 0.077 | -0.356*** | -0.110 | -0.025 |